OpenStudy (anonymous):

What is the sum of the measures of the interior angles of a regular polygon where a single exterior angle measures 72°?

OpenStudy (anonymous):

360/n = 72 360 = 72n n = 5 So it's a pentagon Now use the formula to find the sum of the measures of interior angles

OpenStudy (anonymous):

so its 5 ?

OpenStudy (anonymous):

nope it's a five-side shape

OpenStudy (anonymous):

ok so its 108 ?

OpenStudy (anonymous):

it says sum of all interior angles

OpenStudy (anonymous):

180 * (n-2) is the general formula 180 * (5 - 2) = 180 * 3 = 540

OpenStudy (anonymous):

ok so 5 * 72?

OpenStudy (anonymous):

actually 5 * 108

OpenStudy (anonymous):

right so i was on the right track

OpenStudy (anonymous):

Good Job!

OpenStudy (anonymous):

540

OpenStudy (anonymous):

yes